a solid sphere is rolling down an inclined plane of inclination 30 degrees without slipping the friction force acting on the sphere is
 

Grade:11

2 Answers

siddharth swami
6 years ago
riction oce would be acting in the diretion up the plane. So equation wold be
   mg*sin30 – f = ma   
   f*r = I (alpha)        [I= moment o inertia, alpha= angular acceleration]
alpha = a/r
solving these eqation we get, f = mg/7
   
siddharth swami
6 years ago
 
 
 
 
 
 
 
 
friction would be acting in the diretion up the plane. So equation wold be
   mg*sin30 – f = ma   
   f*r = I (alpha)        [I= moment o inertia, alpha= angular acceleration]
alpha = a/r
solving these eqation we get, f = mg/7